OpenStudy (anonymous):

What is 67% of 22? Round to the nearest whole number if necessary

OpenStudy (anonymous):

67% is 67/100 to find the percentage of something, you multiply it by the equivalent fraction.

OpenStudy (anonymous):

so you would do 67% x 22?

OpenStudy (anonymous):

No. You're trying to multiply 22 by 67% You have to multiply 22 by the fraction EQUIVALENT to 67% Also, if you prefer, ANY number divided by "100" may also be expressed as a decimal: 1% = 1/100 or .01 6% = 6/100 which is also .06 25% = 25/100 which is also .25 67%=67/100 which is also .67 99%=99/100 which is also .99 100%=100/100 which is also 1

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np